OpenStudy (zzr0ck3r):

\((-1) = (-1)^1=(-1)^{\frac{1}{1}}=(-1)^{\frac{2}{2}}=\sqrt{(-1)^2}=|-1|=1\) Go!

OpenStudy (zzr0ck3r):

an openstudy oldie but goodie

OpenStudy (anonymous):

so wat's da question?

OpenStudy (zzr0ck3r):

do you see a problem with what I did?

OpenStudy (anonymous):

theres nothing wrong with that

OpenStudy (zzr0ck3r):

so -1 = 1?

OpenStudy (zzr0ck3r):

god I hope something is wrong, or I just broke all of math

OpenStudy (kainui):

Thanks, now let's see what fun consequences we can derive from this awesome result. 0=1-1=1+1 0=2 Yay

OpenStudy (zzr0ck3r):

lol

OpenStudy (zzr0ck3r):

thus 1/2 = \(\infty\)

OpenStudy (anonymous):

It's hard to say whether \[ (-1)^{\frac 11} = (-1)^{\frac 22} \]or \[ (-1)^{\frac 22} = \sqrt{ (-1)^2} \]is the problem.

OpenStudy (kainui):

By induction, since 0=2, and 2=2+0, then...

OpenStudy (zzr0ck3r):

I think its the first step

OpenStudy (anonymous):

I believe zepp had asked this question a few years ago.

OpenStudy (ikram002p):

\( (-1)^\frac{1}{1}\neq (-1)^\frac{2}{2} \) we should have respect to even odd :P

OpenStudy (zzr0ck3r):

fractions mean something different when raised to an exponent

OpenStudy (zzr0ck3r):

the equivalence is changed in this meaning

OpenStudy (anonymous):

Yeah, and you had answered it lol http://openstudy.com/updates/5014e187e4b0fa2467312775

OpenStudy (kainui):

\[\large -1 = e^{i \pi}\]\[\large \sqrt{e^{i 2\pi}}= \pm 1\]

OpenStudy (zzr0ck3r):

so what you are saying then @ikram002p is that \(1=\frac{2}{2}\ne \frac{1}{1}=1\)?

OpenStudy (ikram002p):

hehe well consedering the power which have its own properties

OpenStudy (zzr0ck3r):

I dont really know the answer, and when I asked my professor for a good explination he laughed at me. I am not quite sure what to take away from that

OpenStudy (anonymous):

The fact that this hurts my head reading it means I probably should get some sleep soon...

OpenStudy (kainui):

This has the whole problem of the fact that any time you take the root we cut it into two parts and just take the real positive part when taking the square root. Damn principal square root. It's like the 4th root could really be a 1-to-4 mapping haha.

OpenStudy (zzr0ck3r):

its good stuff:)

OpenStudy (anonymous):

Perhaps fractional exponents are just bad practice?

OpenStudy (kainui):

\[\sqrt[4]{16}=2=-2=2i=-2i\] The world is broked

OpenStudy (zzr0ck3r):

I think so @wio its bad notation that we need. The better of two devils i suppose.

OpenStudy (anonymous):

i love u, kainui

OpenStudy (kainui):

I guess I have to give you a medal now.

OpenStudy (zzr0ck3r):

another good one. how is it that addition is commutative yet we get a different infinite sum when we move around terms in many infinite sums...

OpenStudy (anonymous):

Consider \[ (-1)^{3/2} \]We have \[ ((-1)^{3})^{1/2} = (-1)^{1/2} = i \]and \[ ((-1)^{1/2})^{3} = (i)^{3} = -i \]

OpenStudy (anonymous):

Oh ya, that classic one.\[S = 1 - 1 + 1 -1 \cdots\]

OpenStudy (zzr0ck3r):

right...

OpenStudy (kainui):

\[\Large \sqrt[1]{-1}=-1\]??? @zzr0ck3r I think that's because those aren't convergent sums.

OpenStudy (zzr0ck3r):

but even some when we change finite amount of terms will change the outcome...

OpenStudy (anonymous):

The problem is that order of operations MATTER and we are trying to suggest that they somehow don't with such notation.

OpenStudy (zzr0ck3r):

even on converging sums

OpenStudy (kainui):

Give an example of a convergent infinite sum that doesn't work. I'm curious, this is my groove.

OpenStudy (zzr0ck3r):

I thought, maybe I am remembering something wrong.

OpenStudy (zzr0ck3r):

sec ill look

OpenStudy (anonymous):

for example \[ (-1) = (\sqrt{-1})^2 = -1 \]But \[ (-1) = \sqrt{-1^2} = 1 \]

OpenStudy (zzr0ck3r):

@Kainui There is a famous theorem of Riemann that says that for a conditionally convergent series, we can re-order the terms so that the series converges to any limit or even diverges.

OpenStudy (zzr0ck3r):

this is not my bag, I am an algebraist, but its still fascinating

OpenStudy (kainui):

Well I guess there is that thing I heard of, where if you have a sphere and cut it up into a bunch of pieces then you can get two spheres of exactly the same size; which seems ridiculous and related.

OpenStudy (zzr0ck3r):

yeah CRAZY!

OpenStudy (anonymous):

Banach-Tarski?

OpenStudy (zzr0ck3r):

I cant wait till topology next term!

OpenStudy (kainui):

Did you see that thing I was talking about a month ago with that geometric series where it converged to the average?

OpenStudy (zzr0ck3r):

go read on the Cantor set/function...

OpenStudy (zzr0ck3r):

a continuous step function....

OpenStudy (zzr0ck3r):

ffs what is going on !!!!!!!!!!!!!!!!!

OpenStudy (zzr0ck3r):

lol

OpenStudy (kainui):

Yeah that sounds exciting, there's a lot of interesting stuff in math that I'll have to get around to teaching myself since I won't be taking it haha. What are either of those?

OpenStudy (zzr0ck3r):

its hard to explain but pretty easy to read

OpenStudy (zzr0ck3r):

actually pretty good sumation of the cool properties here http://en.wikipedia.org/wiki/Cantor_function

OpenStudy (kainui):

Well I have some suspicions about how topology can apply to organic chemistry and knots that I'd like to understand more.

OpenStudy (zzr0ck3r):

sweet, you should explore that. what a great MS research topic for someone in both fields

OpenStudy (zzr0ck3r):

when my professor did his lecture on the Cantor set, he said "today we step into the dark side" (the same guy that laughed at me for the question in the topic here)

OpenStudy (kainui):

I mean, knots and molecules are both three dimensional structures that we try to represent in two dimensions. I have a feeling that some organic synthesis problems might be worth trying to decompose like "skein relations" in knot theory. But I don't really know enough about that sort of stuff or group theory.

OpenStudy (zzr0ck3r):

sounds believable to me!

OpenStudy (kainui):

Hahaha why did he say that?

OpenStudy (zzr0ck3r):

because its very counter intuitive

OpenStudy (zzr0ck3r):

you take a set, and keep removing things from it, and in the end the set has the same length as when you started...

OpenStudy (zzr0ck3r):

by length i mean measure...

OpenStudy (anonymous):

all of dis is interesting. but its still theoretical.

OpenStudy (kainui):

I think if money was not an issue, I'd try to really understand the structure of DNA and cells and try to find the algorithm for how life unfolds. We're symmetrical. We're all "even" functions in a sense haha.

OpenStudy (kainui):

Wait, do you mean like because the set is infinite? For instance I think Galileo said something like there are an infinite number of integers and an infinite number of squares. But it seems like there are fewer squares. But both are infinite, which is interesting.

OpenStudy (zzr0ck3r):

@kainui read that link, its to much to explain, and I would have to reference that anyway...

OpenStudy (anonymous):

...wow.

OpenStudy (zzr0ck3r):

I will be taking my medals after you watch that(if you dont die from laughter).

OpenStudy (kainui):

Lol mean jerk time.

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

i... no comments.

OpenStudy (kainui):

This reminds me of when I was teaching my friend mechanics and was telling him that you can find the initial velocity of... if you time your squirt and measure the distance and height lol.

OpenStudy (anonymous):

is it abnormal to puke and laugh at the same time?

OpenStudy (kainui):

What's Silicon Valley I feel like I would love this movie/show.

OpenStudy (zzr0ck3r):

its on hbo, its about a bunch of coders creating a startup business and getting funding

OpenStudy (anonymous):

Um, not going to acknowledge the fact that: \[ (\sqrt{x})^n \neq \sqrt{x^n} \]And thus \(x^{n/2}\) because it ignores this distinction?

OpenStudy (anonymous):

This is why I originally said what I said before. Either we impose the rule \[ x^{n/2} = (\sqrt{x})^2 \]And the second step is wrong, otherwise \[ x^{1/1} =x^{2/2} \]must be wrong.

OpenStudy (zzr0ck3r):

yeah I always figured it was that step that was the problem because there are counter examples to that claim, I just never had an "axiomatic" reason for it.

OpenStudy (anonymous):

Wolfram actually imposes that rule: http://www.wolframalpha.com/input/?i=%28-1%29%5E%282%2F2%29

OpenStudy (kainui):

@wio yeah, when I think of numbers I basically see every number as being of the form: \[\Large re^{i \theta}\] So you have an angular part and a scalar part. So this means, -1, i, -1, and 1 are really like the "sign" part of your number and can be considered separately. So if you look at -1 as being rotated by 180 degrees then you can interpret the powers as being rotation. Then the question becomes, are you counter rotating as much as you're rotating or have you tricked yourself?

OpenStudy (anonymous):

If you impose that rule, then the notation has meaning. Otherwise the notation itself is unacceptable.

OpenStudy (zzr0ck3r):

But god, can you imagine if we tried to teach that when we learn things like \(\sqrt[a]{b^c}=b^{\frac{c}{a}}\).

OpenStudy (zzr0ck3r):

this is my favorite part about this site...

OpenStudy (anonymous):

That rule is false when \(a\equiv 0\pmod 2\). It's true only half the time and it gets taught. It really needs to be fixed.

OpenStudy (anonymous):

Or actually, the problem of that rule is only when \(b\) is negative.

OpenStudy (zzr0ck3r):

I think if we only reduce/scale fractions when there denominator is coprime with the numerator we are ok.

OpenStudy (zzr0ck3r):

when in lowest terms....

OpenStudy (zzr0ck3r):

wait

OpenStudy (zzr0ck3r):

that makes no sense....

OpenStudy (anonymous):

You don't want 2s canceling?

OpenStudy (zzr0ck3r):

right

OpenStudy (zzr0ck3r):

2/4 cant go to 1/2 but 3/6 can, I think

OpenStudy (anonymous):

I think rule should be \[ \sqrt[a]{x^b} = (x^b)^{1/a} \]Then it should be taught \[ (x^b)^c = x^{bc} \iff x\geq 0 \]